# On the Robustness Tradeoff in Fine-Tuning

### v1

@@ 4.4 On Out-of-Distribution Robustness
Similarly, we track OOD robustness throughout fine-tuning and compare it to in-domain (test and training) accuracy, as shown in Figure 6. In comparison to adversarial robustness, which often deteriorates after peaking, OOD robustness plateaus after the initial improvement at a lower level relative to standard accuracy. Both in-domain and OOD robustness slightly decline after convergence, likely due to traditional overfitting rather than the accuracy-robustness conflict in adversarial settings. This behavior highlights the different mechanisms behind adversarial and OOD robustness. While adversarial robustness is affected by low-level, non-robust features, OOD robustness depends on transferable features that are less sensitive to fine-tuning-induced degradation.
As shown in Figure 7, we further analyze OOD robustness across the seven fine-tuning strategies and six training domains. Linear probing consistently yields the lowest OOD robustness ($61\%\pm 5\%$), while full fine-tuning achieves the highest ($73\%\pm 2\%$). In addition, models fine-tuned on the “real” domain, which is closest to the pre-training distribution, exhibit lower OOD robustness ($64\%\pm 5\%$) compared to more shifted domains such as “infograph” ($73\%\pm 4\%$) and “quickdraw” ($72\%\pm 3\%$). Interestingly, OOD robustness stays stable across fine-tuning strategies, excluding linear probing. This suggests that fine-tuning methods supports model generalization to other domains similarly. The results reinforce that OOD robustness is primarily dependent on domain shifts and the extent of parameter updates (e.g. linear probing vs. full fine-tuning), rather than the specific underlying mechanisms.
@@ 5 Related Work
Furthermore, with the rise of fine-tuning, robustness considerations of them have become increasingly important. Prior research has explored robustness from a broader perspective, focusing on studying adversarial robustness either during (a) pre-training robustness [4, 21, 6] or (b) fine-tuning [47, 34, 20]. Specifically, these studies primarily investigate full fine-tuning or linear probing, without considering various PEFT strategies. Recent studies have attempted to improve PEFT robustness. Hua et al. [17] propose a robustness-aware initialization strategy. Nguyen and Le [38] integrate adversarial training with adapters using mixup. However, they focus on the final model state instead of investigating with a finer granularity on how the robustness-accuracy tension change throughout fine-tuning and its implications.
6 Discussion
As adversarial robustness is closely tied to non-robust features from downstream data, one natural question arises: are there non-robust features from upstream data? The “robustness” of features here is relative. Since the goal is to apply models to solve downstream tasks, we only consider model accuracy and robustness on downstream phenomena. Thus, non-robust features specific to upstream data may be robust to adversarial perturbations targeted at downstream data. However, it is convoluted to distinguish when the model learns what features from different data phenomena when models can also learn them simultaneously. One potential way to further investigate this is to track robustness using adversarial examples generated on both upstream and downstream data.
@@ 7 Conclusion
Our study systematically examines the robustness-accuracy trade-off in the fine-tuning paradigm, revealing that adversarial robustness initially improves but declines as fine-tuning continues. Across 231 fine-tuned models with 7 state-of-the-art fine-tuning strategies and our independent security and safety evaluation framework, we find that the balance of the trade-off varies with fine-tuning strategies, downstream data complexity, and its similarity to pre-training data. Specifically, methods updating attention-related layers (e.g., LoRA, Compacter) tend to better balance robustness and accuracy, while simpler adaptation techniques (e.g., BitFit) achieve higher robustness peaks but degrade faster. However, this trade-off does not extend to safety (OOD) robustness, which instead depends on the model’s ability to generalize across domains. These findings suggest that the design of robustness aware fine-tuning strategies should consider both adversarial and OOD robustness independently.
In the future, we plan to extend our evaluation framework to defended models (e.g., with adversarial training) and other security (e.g., black-box, adaptive attacks) and safety (e.g., corrupted data) environments.
@@ A.2 Trade-off Space of Fine-Tuning
A.3 Hyperparameters
@@
A.4 Decomposition of Fine-Tuning
As described in Section 3.2.1, we focus on decomposing PEFT methods here along two directions: information location and mechanisms, each having four components. The decomposition for five PEFT methods can be found in Table 5.

### v2

@@ 4.4 On Out-of-Distribution Robustness
Similarly, we track OOD robustness during fine-tuning and compare it to in-domain (test and training) accuracy, as shown in Figure 6. In comparison to adversarial robustness, which often deteriorates after peaking, OOD robustness plateaus after the initial improvement at a lower level than standard accuracy. Both in-domain and OOD robustness slightly decline after convergence, likely due to overfitting rather than the accuracy-robustness conflict in adversarial settings. This behavior highlights the different mechanisms behind adversarial and OOD robustness. While adversarial robustness is affected by low-level, non-robust features, OOD robustness depends on transferable features that are less sensitive to fine-tuning-induced degradation.
As shown in Figure 7, we further analyze OOD robustness across the seven fine-tuning strategies and six training domains. Linear probing consistently yields the lowest OOD robustness ($61\%\pm 5\%$), while full fine-tuning achieves the highest ($73\%\pm 2\%$). In addition, models fine-tuned on the “real" domain, which is closest to the pre-training distribution, exhibit lower OOD robustness ($64\%\pm 5\%$) compared to more shifted domains such as “infograph" ($73\%\pm 4\%$) and “quickdraw" ($72\%\pm 3\%$). Interestingly, OOD robustness stays stable across fine-tuning strategies, excluding linear probing. This suggests that fine-tuning methods support model generalization to other domains similarly. The results reinforce that OOD robustness is primarily dependent on domain shifts and the extent of parameter updates (e.g. linear probing vs. full fine-tuning), rather than the specific underlying mechanisms.
@@ 5 Related Work
Furthermore, with the rise of fine-tuning, robustness considerations of them have become increasingly important. Prior research has explored robustness from a broader perspective, focusing on studying adversarial robustness either during (a) pre-training robustness [4, 22, 6] or (b) fine-tuning [48, 35, 21]. Specifically, these studies primarily investigate full fine-tuning or linear probing, without considering various PEFT strategies. Recent studies have attempted to improve PEFT robustness. Hua et al. [18] propose a robustness-aware initialization strategy. AdapterMixup [39] integrates adversarial training with adapters using mixup. CLAT [12] studies layer-level robustness under adversarial training, while we focus on method-level evaluation across fine-tuning strategies. Additionally, they study the final model state instead of investigating with a finer granularity on how the robustness-accuracy tension changes fundamentally throughout fine-tuning before introducing robustness-improving mechanisms.
6 Discussion & Limitations
Non-robust features & Robustness. As adversarial robustness is closely tied to non-robust features from downstream data, one natural question arises: are there non-robust features from upstream data? The “robustness" of features here is relative. Since the goal is to apply models to solve downstream tasks, we only consider model accuracy and robustness on downstream phenomena. Thus, non-robust features specific to upstream data may be robust to adversarial perturbations targeted at downstream data. However, it is convoluted to distinguish when the model learns what features from different data phenomena when models can also learn them simultaneously. One potential way to further investigate this is to track robustness using adversarial examples generated on both upstream and downstream data.
Limitations & future work. Our study focuses on understanding robustness dynamics of SOTA PEFT strategies before applying defense mechanisms. Due to runtime constraints in continuous evaluation, we use PGD, which has been shown to reliably estimate robustness for clean models [7]. Future work can extend our framework to additional model types (e.g., CNN-ViT hybrid models, adversarially-trained models) and broader threat scenarios (e.g., black-box, adaptive, and data corruption attacks).
@@ 7 Conclusion
Our study systematically examines the trade-off between robustness and accuracy in the fine-tuning paradigm, revealing that adversarial robustness initially improves but declines as fine-tuning continues. Across 231 fine-tuned models with 7 state-of-the-art fine-tuning strategies and our independent security and safety evaluation framework, we find that the balance of the trade-off varies with fine-tuning strategies, downstream data complexity, and its similarity to pre-training data. Specifically, methods updating attention-related layers (e.g., LoRA, Compacter) tend to better balance robustness and accuracy, while simpler adaptation techniques (e.g., BitFit) achieve higher robustness peaks but degrade faster. However, this trade-off does not extend to safety (OOD) robustness, which instead depends on the model’s ability to generalize across domains. These findings suggest that the design of robustness-aware fine-tuning strategies should consider both adversarial and OOD robustness independently.
Acknowledgements. We would like to thank Kassem Fawaz and Rahul Chatterjee for their helpful comments on earlier iterations of this work. This material is based upon work supported by the National Science Foundation under Grant No. CNS-2343611 and by the U.S. Army Research Office under MURI grant No. W911NF-21-1-0317. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the U.S. Army Research Office. The U.S. Government is authorized to reproduce and distribute reprints for government purposes notwithstanding any copyright notation hereon.
@@ A.2 Trade-off Space of Fine-Tuning
A.3 Ablation Study
We perform an ablation study on the number and location of trainable parameters and learning rate. We conduct experiment sweeps with final model checkpoints on 1) LoRA rank (r $\in$ [1,20]), 2) Adapter reduction factor ($d\in$ [4,32]), and 3) update location (attention vs. FNN). The results show that robustness does not consistently correlate with the number of trainable parameters or layers updated. This suggests that these factors alone do not explain robustness differences. However, the results all show low adversarial accuracy with small variances. It led us to our design choise: tracking changes over training steps offer deeper insights.
In addition, we also study how different learning rates affect OOD robustness during training. We track OOD robustness while fine-tuning models with varying learning rates—$\{1e-4,5e-4,1e-3,5e-3\}$. As shown in Figure 9, learning rate does impact accuracy and robustness, especially during the early steps, but show consistent trend as claimed across steps and converge in the end.
Figure 9: The OOD robustness of models fine-tuned by LoRA with different learning rates on the “real" domain and evaluated on 5 other domains.
A.4 Hyperparameters
@@
A.5 Dataset Scale
Figure 10: Continuous evaluation of training accuracy (blue), test accuracy (green), and adversarial robustness (red) across backpropagation steps (truncated at 10,000 steps) on Places365 [52].
PEFT is especially relevant in low-data regimes. Our main experiments focus on datasets with 10k to 60k samples, covering diverse degrees of task complexity in terms of number of classes, class separation, and similarity to upstream data. To complement this analysis and assess the effect of data scale, we also include experiments on Places365 [52], a medium-scale dataset with approximately 1.8 million samples. As shown in Figure 10, models achieve relatively low clean accuracy (preaking around 50%) and limited adversarial robustness (less than $5\%$). Due to this overall weak performance, the trend of adversarial robustness and the differences of robustness across fine-tuning strategies are difficult to distinguish.
A.6 Decomposition of Fine-Tuning
As described in Section 3.2.1, we focus on decomposing PEFT methods along two directions: information location and mechanisms, each having four components. The decomposition for five PEFT methods can be found in Table 5.